trevor bought a diamond pendant with a purchase price of 363.when sales tax was added ,he paid an overall price of 392.04.what was the tax rate
Answer by ewatrrr(24785)

363 + 363r = 392.04
363r = 29.04
r = 29.04/363 = .08 0r 8%